------------------------------------------------------------ Date: Mon Jul 3 00:01:10 2000, Shift: 16:00-24:00 Supervisor: KO, with: BS, PC Log book#2 pages 118 - 118 Shift Summary: ------------------------------------------------------------ 1. Machine status: up all shift, cycling through "inject, ramp, drop the beam" cycles about every 40 minutes. Had one 2 hour store between 21:00 and 23:00 with lousy intensity, blue starting at 250 dropping to 150, yellow steady at 400. A good loocking store starting now at 23:45 with blue at 400 and yellow at 240. 2. Took pedestal runs and a data run during the 21:00 to 23:00 store with lousey intensity. 3. Problems discovered during shift check: 1) three bad HV cards (two on FFS L1440 (H1 and C1), one in BFS L1440 (H2)). These will be replaced during downtime later this week. 2) we need better "summary" histograms in SuperMon for each detector. Ideally there should be as few as one or two histograms per detector, for example showing the ADC mean and width versus channel number. The current scheme where one has to look at a gazillion histograms displayed with an unreadable font is useless. 3) in the FEH, while checking the power-up status of bins and crates, it was found that at least 50% of all bins have burned-out power lights. These gotta be fixed. 4) in the gas room, it is not clear what we should look for when checking the gas cylinders. Maybe the pressure of each cylinder should be recorded in the checklist? 5) the control room was found in an unclean state: scrap paper everywhere, unwashed coffe mugs, no water, no sugar. Please keep the control room clean- dispose of all paper into the recycle bin, wash the cups after use, and bring more water and sugar from Bldg.510! K.O. ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Mon Jul 03 2000 - 00:05:20 EDT